if tan2x.tan7x=1.Then what is valueof tan3x

Grade:12th Pass

2 Answers

Rohan Das
11 years ago

tan9x=tan(7x+2x)=(tan7x+tan2x)/(1-tan7xtan2x)

but given tan7xtan2x=1

so tan (7x+2x)=(tan7x+tan2x)/(1-tan7xtan2x)

that is =(tan7x+tan2x)/(1-1)

which is not defined.

so 9x=7x+2x=90degree

hence 9x=90degree

so 3x=30degree

Akash Kumar Dutta
11 years ago

1/root(3)....[ x=10 degree ]